Find the area of the region inside the cardioid r= 4-4sintheta and outside the the circle r=6.
Answer #1

4-4 sino = 6 -4 sine = 2 =) sino - -Ź 0 -- and @ = 21 + (916-16-16viño su sinode - t o - tosiño saudaro de = + (20-10 (1 ) as